# Expanding the Chaos: Neural Operator for Stochastic (Partial) Differential Equations

### v1

@@ F.4 Experimental Details for TSBM
Similar to the motivation of $\mathcal{U}$-SDENO, we design the $\mathcal{G}$-SDENO followed by the fact that graph neural networks [kipf2016semi] can be served as the feature propagator of the brain regions. Different from the $\mathcal{U}$-SDENO, in GNNs, there is no time encoding inside; accordingly, time embedding is conducted outside the GNN process. Two outpus, i.e., GNN feature and time embedding, will be merged by inner product.
@@
The brain graph data used in this study is constructed sourced from the Human Connectome Project (HCP) Young Adult dataset [van2013wu]. The graph topology is defined by the Glasser multi-modal parcellation (MMP1.0), which delineates 360 distinct cortical areas that serve as the nodes of our graph [glasser2016multi].
To determine the spatial embedding for each of the 360 nodes, we processed the HCP’s standard 32k fs_LR cortical surface meshes. First, for each cerebral hemisphere, we identified the complete set of vertices belonging to each of the 360 parcels using the provided .dlabel file. The definitive 3D coordinate for each node (parcel) was then computed as the median of the coordinates of all its constituent vertices. This robust aggregation provides a geometrically central and stable position for each brain region. The edges of the graph, representing the underlying functional connectivity, are determined from a sparse inverse covariance matrix estimated from the fMRI signals, following the methodology in yang2025topological. The features associated with each node are derived from task-based functional MRI (tfMRI) signals corresponding to a working memory task. Specifically, our experiment aims to model the transition of brain states from a high-energy "2-back" working memory task (the initial distribution) to a lower-energy "0-back" task (the target distribution), providing a meaningful context for evaluating the topological interpolation capabilities of the model [yang2025topological].
@@
The pretraining of the Topological Schrödinger Bridge Matching (TSBM) model follows the alternating training scheme. The process involves iteratively training a forward SDE and a backward SDE to progressively match the initial and target distributions. In our brain signal experiment, this corresponds to mapping the "2-back" task brain states to the "0-back" task states. Each SDE is driven by a GNN with 4 residual blocks, tasked with learning the drift term of the process. We employ a variance-exploding (VE) SDE formulation with a noise parameter of $\sigma=0.5$. The model is trained for a total of 2000 iterations. For each iteration in the alternating loop, we first train the forward SDE for one epoch and then train the backward SDE for one epoch. The training is performed using the Adam optimizer with a learning rate of $10^{-3}$ for both networks and a batch size of 64. A step-wise learning rate scheduler is used, reducing the learning rate by a factor of 0.5 every 500 iterations to ensure stable convergence. Finally, the model is evaluated via the 1-Wasserstein distance, resulting in 9.51.
@@
The training of $\mathcal{G}$-SDENO is similar to $\mathcal{U}$-SDENO, except the time embedding is conducted outside the GNN propagation; therefore, we omit it here. One interesting observation we found in our experiment is that our results are generally invariant across the GNNs, such as GCN [kipf2016semi], and ChebyNet [defferrard2016convolutional]. We initially believe that this is due to the fact that all these GNNs will tend to smooth the node features so that the feature difference will asymptotically become identical, resulted in the over-smoothing problem (OSM). To this end, we test those GNNs that can provably avoid the OSM problem, such as generalised framelet model [han2022generalized]. To our surprise, even when the GCN model is high-frequency dominant, meaning that node features tend to distinguish from each other through the propagation, we will find that both TSBM and our model show similar results compared to those GCN and ChebyNet. That may lead to a deeper exploration on the relationship between the diffusion task (i.e., interpolation) and the denoising models (propagator networks in our case), we leave this task to future works.
@@ F.6 Experimental Details for Parameter Estimation
where $\delta_{\boldsymbol{\gamma}}$ is the Dirac measure at position $\boldsymbol{\gamma}$. As established in the foundational work on this topic, in the mean-field limit where the number of ensemble members $\ell\to\infty$, this empirical measure converges weakly to a deterministic probability measure $\mu_{t}$ [law2016deterministic, ertel2025mean]. The evolution of this limiting measure is described by a non-linear Fokker-Planck type partial differential equation, often referred to as a McKean-Vlasov equation [chan1994dynamics].
@@ F.7 WCE on Manifold
In this section, we show experimental details on our flood prediction, which employs the SDE on the sphere. Similarly, we pretrain the Riemannian score based generative models [de2022riemannian] on the Flood dataset, and sample 200 Riemannian Brownian increments on the manifold tangent bundle.
@@
Manifold Constraints. Third, our manifold SDE experiment on the sphere $\mathbb{S}^{2}$ relies on a practical simplification: we linearize the problem by parallel-transporting the entire stochastic process to a single fixed tangent plane and constructing Wick features there. This “global flattening” works well on the sphere, a manifold with constant positive curvature, but its generalization to more complex Riemannian manifolds is non-trivial. On manifolds with varying curvature or complicated topology, a single global reference frame may not exist, and parallel transport can be path-dependent and introduce distortion, potentially degrading the statistical properties of the transported Brownian motion. Intrinsic manifold-native constructions, such as Wiener–chaos expansions built directly from the isometry group on specific spaces [kalpinelli2013numerical], avoid these issues but require sophisticated, manifold-dependent machinery. Extending SDENO to such intrinsic constructions, or designing robust local/patch-wise linearization schemes on general manifolds, is left for future work.
Singular SPDEs. Finally, our empirical evaluation focuses on non-singular SPDEs (e.g., dynamic $\boldsymbol{\Phi}^{4}_{1}$ in 1D and stochastic Navier–Stokes at moderate resolutions) and several downstream SDE-based tasks (image diffusion, brain graphs, and manifold flows). We do not yet evaluate SDENO on truly singular SPDEs that require renormalisation, where regularity-structure-based neural operators [hu2022neural, gong2023deep] are particularly well suited. A natural next step is to combine our chaos-based noise representation with such advanced architectures and to benchmark SDENO on more challenging singular SPDE datasets, for example those in recent SPDE benchmarks.

### v2

@@ F.4 Experimental Details for TSBM
Similar to the motivation of $\mathcal{U}$-SDENO, we design the $\mathcal{G}$-SDENO followed by the fact that graph neural networks (Kipf and Welling, 2017) can be served as the feature propagator of the brain regions. Different from the $\mathcal{U}$-SDENO, in GNNs, there is no time encoding inside; accordingly, time embedding is conducted outside the GNN process. Two outpus, i.e., GNN feature and time embedding, will be merged by inner product.
@@
The brain graph data used in this study is constructed sourced from the Human Connectome Project (HCP) Young Adult dataset (Van Essen et al., 2013). The graph topology is defined by the Glasser multi-modal parcellation (MMP1.0), which delineates 360 distinct cortical areas that serve as the nodes of our graph (Glasser et al., 2016).
To determine the spatial embedding for each of the 360 nodes, we processed the HCP’s standard 32k fs_LR cortical surface meshes. First, for each cerebral hemisphere, we identified the complete set of vertices belonging to each of the 360 parcels using the provided .dlabel file. The definitive 3D coordinate for each node (parcel) was then computed as the median of the coordinates of all its constituent vertices. This robust aggregation provides a geometrically central and stable position for each brain region. The edges of the graph, representing the underlying functional connectivity, are determined from a sparse inverse covariance matrix estimated from the fMRI signals, following the methodology in Yang (2025). The features associated with each node are derived from task-based functional MRI (tfMRI) signals corresponding to a working memory task. Specifically, our experiment aims to model the transition of brain states from a high-energy ”2-back” working memory task (the initial distribution) to a lower-energy ”0-back” task (the target distribution), providing a meaningful context for evaluating the topological interpolation capabilities of the model (Yang, 2025).
@@
The pretraining of the Topological Schrödinger Bridge Matching (TSBM) model follows the alternating training scheme. The process involves iteratively training a forward SDE and a backward SDE to progressively match the initial and target distributions. In our brain signal experiment, this corresponds to mapping the ”2-back” task brain states to the ”0-back” task states. Each SDE is driven by a GNN with 4 residual blocks, tasked with learning the drift term of the process. We employ a variance-exploding (VE) SDE formulation with a noise parameter of $\sigma=0.5$. The model is trained for a total of 2000 iterations. For each iteration in the alternating loop, we first train the forward SDE for one epoch and then train the backward SDE for one epoch. The training is performed using the Adam optimizer with a learning rate of $10^{-3}$ for both networks and a batch size of 64. A step-wise learning rate scheduler is used, reducing the learning rate by a factor of 0.5 every 500 iterations to ensure stable convergence. Finally, the model is evaluated via the 1-Wasserstein distance, resulting in 9.51.
@@
The training of $\mathcal{G}$-SDENO is similar to $\mathcal{U}$-SDENO, except the time embedding is conducted outside the GNN propagation; therefore, we omit it here. One interesting observation we found in our experiment is that our results are generally invariant across the GNNs, such as GCN (Kipf and Welling, 2017), and ChebyNet (Defferrard et al., 2016). We initially believe that this is due to the fact that all these GNNs will tend to smooth the node features so that the feature difference will asymptotically become identical, resulted in the over-smoothing problem (OSM). To this end, we test those GNNs that can provably avoid the OSM problem, such as generalised framelet model (Han et al., 2022). To our surprise, even when the GCN model is high-frequency dominant, meaning that node features tend to distinguish from each other through the propagation, we will find that both TSBM and our model show similar results compared to those GCN and ChebyNet. That may lead to a deeper exploration on the relationship between the diffusion task (i.e., interpolation) and the denoising models (propagator networks in our case), we leave this task to future works.
@@ F.6 Experimental Details for Parameter Estimation
where $\delta_{\boldsymbol{\gamma}}$ is the Dirac measure at position $\boldsymbol{\gamma}$. As established in the foundational work on this topic, in the mean-field limit where the number of ensemble members $\ell\to\infty$, this empirical measure converges weakly to a deterministic probability measure $\mu_{t}$ (Law et al., 2016; Ertel, 2025). The evolution of this limiting measure is described by a non-linear Fokker-Planck type partial differential equation, often referred to as a McKean-Vlasov equation (Chan, 1994).
@@ F.7 WCE on Manifold
In this section, we show experimental details on our flood prediction, which employs the SDE on the sphere. Similarly, we pretrain the Riemannian score based generative models (De Bortoli et al., 2022) on the Flood dataset, and sample 200 Riemannian Brownian increments on the manifold tangent bundle.
@@
Manifold Constraints. Third, our manifold SDE experiment on the sphere $\mathbb{S}^{2}$ relies on a practical simplification: we linearize the problem by parallel-transporting the entire stochastic process to a single fixed tangent plane and constructing Wick features there. This “global flattening” works well on the sphere, a manifold with constant positive curvature, but its generalization to more complex Riemannian manifolds is non-trivial. On manifolds with varying curvature or complicated topology, a single global reference frame may not exist, and parallel transport can be path-dependent and introduce distortion, potentially degrading the statistical properties of the transported Brownian motion. Intrinsic manifold-native constructions, such as Wiener–chaos expansions built directly from the isometry group on specific spaces (Kalpinelli et al., 2013), avoid these issues but require sophisticated, manifold-dependent machinery. Extending SDENO to such intrinsic constructions, or designing robust local/patch-wise linearization schemes on general manifolds, is left for future work.
Singular SPDEs. Finally, our empirical evaluation focuses on non-singular SPDEs (e.g., dynamic $\boldsymbol{\Phi}^{4}_{1}$ in 1D and stochastic Navier–Stokes at moderate resolutions) and several downstream SDE-based tasks (image diffusion, brain graphs, and manifold flows). We do not yet evaluate SDENO on truly singular SPDEs that require renormalisation, where regularity-structure-based neural operators (Hu et al., 2022; Gong et al., 2023) are particularly well suited. A natural next step is to combine our chaos-based noise representation with such advanced architectures and to benchmark SDENO on more challenging singular SPDE datasets, for example those in recent SPDE benchmarks.